# AgentThink: A Unified Framework for Tool-Augmented Chain-of-Thought Reasoning in Vision-Language Models for Autonomous Driving

### v3

@@ Abstract
Vision-Language Models (VLMs) show promise for autonomous driving, yet their struggle with hallucinations, inefficient reasoning, and limited real-world validation hinders accurate perception and robust step-by-step reasoning. To overcome this, we introduce AgentThink, a pioneering unified framework that, for the first time, integrates Chain-of-Thought (CoT) reasoning with dynamic, agent-style tool invocation for autonomous driving tasks. AgentThink’s core innovations include: (i) Structured Data Generation, by establishing an autonomous driving tool library to automatically construct structured, self-verified reasoning data explicitly incorporating tool usage for diverse driving scenarios; (ii) A Two-stage Training Pipeline, employing Supervised Fine-Tuning (SFT) with Group Relative Policy Optimization (GRPO) to equip VLMs with the capability for autonomous tool invocation; and (iii) Agent-style Tool-Usage Evaluation, introducing a novel multi-tool assessment protocol to rigorously evaluate the model’s tool invocation and utilization. Experiments on the DriveLMM-o1 benchmark demonstrate AgentThink significantly boosts overall reasoning scores by 53.91% and enhances answer accuracy by 33.54%, while markedly improving reasoning quality and consistency. Furthermore, ablation studies and robust zero-shot/few-shot generalization experiments across various benchmarks underscore its powerful capabilities. These findings highlight a promising trajectory for developing trustworthy and tool-aware autonomous driving models.
footnotetext: Equal contribution.footnotetext: Corresponding authors.footnotetext: This work was completed during the internships of Kangan Qian and Ziang Luo in Xiaomi Corporation.
@@ 1 Introduction
Recent advances in foundation models have opened new opportunities for autonomous driving, where pretrained Large Language Models (LLMs) [10, 2] and Vision-Language Models (VLMs) [35, 47, 28] are increasingly employed to enable high-level scene understanding, commonsense reasoning, and decision making. These models aim to transcend traditional perception pipelines—which rely on hand-crafted components such as object detection [20, 32], motion prediction [29, 40], and rule-based planning [5]—by providing richer semantic representations and broader generalization grounded in web-scale knowledge.
Many recent approaches recast autonomous driving tasks as visual question answering (VQA) problems, applying supervised fine-tuning (SFT) to foundation VLMs with task-specific prompts for object identification, risk prediction, or motion planning [33, 25, 44, 8, 36]. However, as shown in Fig.2 (a), these models typically treat reasoning as a static input-to-output mapping, neglecting the uncertainty, complexity, and verifiability essential to real-world decision making. Consequently, they often suffer from poor generalization, hallucinated outputs, and limited interpretability [42].
To improve robustness and transparency, recent works have explored incorporating chain-of-thought (CoT) reasoning into VLMs, as shown in Fig. 2 (b). Some approaches adopt rigid CoT templates [35, 14], promoting structured logic at the expense of flexibility. Others use open-ended reasoning formats [26, 15], but may overfit to token patterns and exhibit shallow or redundant reasoning. Moreover, most existing methods rely purely on imitation learning from curated trajectories, lacking the ability to detect knowledge uncertainty or invoke tools for intermediate verification [46].
@@
  • We propose AgentThink, the first framework to integrate dynamic, agent-style tool invocation into vision-language reasoning for autonomous driving tasks.
  • We develop a scalable data generation pipeline that produces structured, self-verified data with integrated tool usage and reasoning chains.
  • We introduce a two-stage training pipeline that combines SFT with GRPO, enabling models to learn when and how to invoke tools to enhance reasoning performance.
  • We design new evaluation metrics tailored to autonomous driving tool invocation, capturing tool selection, integration quality, and reasoning-tool alignment.

### v4

# AgentThink[Uncaptioned image]: A Unified Framework for Tool-Augmented Chain-of-Thought Reasoning in Vision-Language Models for Autonomous Driving

@@ Abstract
Vision-Language Models (VLMs) show promise for autonomous driving, yet their struggle with hallucinations, inefficient reasoning, and limited real-world validation hinders accurate perception and robust step-by-step reasoning. To overcome this, we introduce AgentThink, a pioneering unified framework that, for the first time, integrates Chain-of-Thought (CoT) reasoning with dynamic, agent-style tool invocation for autonomous driving tasks. AgentThink’s core innovations include: (i) Structured Data Generation, by establishing an autonomous driving tool library to automatically construct structured, self-verified reasoning data explicitly incorporating tool usage for diverse driving scenarios; (ii) A Two-stage Training Pipeline, employing Supervised Fine-Tuning (SFT) with Group Relative Policy Optimization (GRPO) to equip VLMs with the capability for autonomous tool invocation; and (iii) Agent-style Tool-Usage Evaluation, introducing a novel multi-tool assessment protocol to rigorously evaluate the model’s tool invocation and utilization. Experiments on the DriveLMM-o1 benchmark demonstrate AgentThink significantly boosts overall reasoning scores by 53.91% and enhances answer accuracy by 33.54%, while markedly improving reasoning quality and consistency. Furthermore, ablation studies and robust zero-shot/few-shot generalization experiments across various benchmarks underscore its powerful capabilities. These findings highlight a promising trajectory for developing trustworthy and tool-aware autonomous driving models. Code is available at https://github.com/curryqka/AgentThink.
@@ 1 Introduction
Recent advances in foundation models have opened new opportunities for autonomous driving [18], where pretrained Large Language Models (LLMs) [10, 2] and Vision-Language Models (VLMs) [37, 49, 28] are increasingly employed to enable high-level scene understanding, commonsense reasoning, and decision making. These models aim to transcend traditional perception pipelines, which rely on hand-crafted components such as object detection [31, 34], motion prediction [30, 42], and rule-based planning [5] by providing richer semantic representations and broader generalization grounded in web-scale knowledge.
Many recent approaches recast autonomous driving tasks as visual question answering (VQA) problems, applying supervised fine-tuning (SFT) to foundation VLMs with task-specific prompts for object identification, risk prediction, or motion planning [35, 25, 46, 8, 38]. This paradigm are widely used in offline pipelines to mine rare driving scenarios, such as identifying intersections with a typical traffic signals or dense pedestrian activities. However, as shown in Fig.2 (a), these models typically treat reasoning as a static input-to-output mapping, neglecting the uncertainty, and verifiability essential to real-world decision making. Consequently, they often suffer from poor generalization, hallucinated outputs, and limited interpretability [44].
@@
To improve robustness and transparency, recent works have explored incorporating chain-of-thought (CoT) reasoning into VLMs, as shown in Fig. 2 (b). Some approaches adopt rigid CoT templates [37, 14], promoting structured logic at the expense of flexibility. Others use open-ended reasoning formats [26, 15], but may overfit to token patterns and exhibit shallow or redundant reasoning. Moreover, most existing methods rely purely on imitation learning from curated trajectories, lacking the ability to detect knowledge uncertainty or invoke tools for intermediate verification [48, 28, 29].
@@
  • Propose AgentThink, the first framework to integrate dynamic, agent-style tool invocation into vision-language reasoning for autonomous driving tasks. It significantly reduces hallucination in VQA and improves overall reasoning consistency by grounding each reasoning step in reliable tool outputs.
  • Develop a scalable data generation pipeline that produces structured, self-verified data with integrated tool usage and reasoning chains.
  • Introduce a two-stage training pipeline that combines SFT with GRPO, enabling models to learn when and how to invoke tools to enhance reasoning performance.
  • Design new evaluation metrics tailored to autonomous driving tool invocation, capturing tool selection, integration quality, and reasoning-tool alignment.